OpenStudy (lena772):

The anion of a Weak Acid is A) a strong base B) a strong acid C) a weak base D) a weak acid E) amphiprotic F) a spectator ion

OpenStudy (lena772):

@hba

OpenStudy (photon336):

Well a weak acid is only partially dissociated, so the anion or conjugate base would be a strong base I think. Remember whenever an acid loses a proton the result is the conjugate base; when a base gains a proton the result is the conjugate acid.

OpenStudy (lena772):

Strong base is incorrect. :/ @Photon336

OpenStudy (lena772):

@Haseeb96 @Zarkon

OpenStudy (aaronq):

Are weak bases. Look at the relationship \(K_w=K_A*K_B\) if we have a weak acid, like HF, with \(K_A=7.2*10^{-4}~M\) then we have \(K_B=\dfrac{K_W}{K_A}=\dfrac{1.0*10^{-14}M^2}{7.2*10^{-4}~M}=1.39*10^{-11}~M\) which is a value expected for a weak base.

OpenStudy (cuanchi):

So @aaronq , doesnt matter if the acid is strong or weak, the conjugated base (the anion of the acid) always will be a weak base? I was also confused by the sentence "the stronger the acid, the weaker the conjugated base" that take me to think that the weakest the acid the strongest the conjugated base as @Photon336 mention. But according to your explanation even the weakest acid will produce a weak conjugated base.

OpenStudy (aaronq):

Yeah. It'll always be weak (relatively, look up the Kb for strong bases and compare), if it wasn't it would immediately mop up the proton and the substance would not be acidic.

OpenStudy (photon336):

@aaronq interesting. point. so it has to be weak because if it immediately pulled up the protons it wouldn't be acidic to begin with. very interesting

OpenStudy (cuanchi):

@aaronq thank you !!!

OpenStudy (aaronq):

no problem !